# Colorful Minors

### v1

@@ 6.1 The sets and
7 A colorful minor algorithm
With lemma 3.9 we have already established our main combinatorial tool for solving Colorful Minor Checking. Indeed, in this section we present an algorithm that solves a common generalization of both Colorful Minor Checking and Weak Disjoint Colorful Paths.
Combining our algorithm for Colorful Minor Checking with theorem 1.1 and revisiting the classic paper of Fellows and Langston [26], implies that for any parameter defined on -colorful graphs and which is monotone under taking colorful minors there exists an FPT-algorithm for deciding
We need to introduce some additional concepts from [79] in order to describe the problem which we want to find an algorithm for.
A rooted (-colorful) graph is a tuple where is a (-colorful) graph and (not necessarily distinct). Isomorphisms between rooted (-colorful) graphs are defined in the natural way. That is and are isomorphic if there exists an isomorphism between and such that for all
We present the following definitions only for rooted colorful graphs, however the definitions for rooted graphs can be derived by omitting all references to colorings.
Let and be rooted -colorful graphs. A model of the second in the first is a minor model of in such that
i)
for all and
ii)
for all
If there is such a model in we say that is a rooted colorful minor of
We call the set of all rooted colorful minors of the folio of Notice that the folio is finite up to isomorphisms. Now let be an integer. We say that has detail at most if and The -folio of is the set of all minors of with detail at most It follows that the size of the -folio of a rooted -colorful graph is always bounded from above by a function of and
Let with The -folio of relative to is the -folio of for any ordering of
The colorful -Folio problem has as input two integers and a -colorful graph and a set
The task is to determine the -folio relative to in
The algorithm for -Folio given by Robertson and Seymour in [79] consists of three different components.
If the input graph contains a large, with respect to and clique-minor, then there exists a vertex whose removal yields and equivalent instance (i.e. the vertex is irrelevant).
So one may assume that the maximum size of a clique minor in is bounded in a function of and If the treewidth of is still large this must mean that there is a large flat wall which again must contain an irrelevant vertex.
Finally, one may assume that the treewidth of is small. Now the problem can be solved by dynamic programming along a tree-decomposition of small width.
It turns out that we only really have to deal with the first of the three steps above by hand. In section 7.2 we provide a reduction of Colorful -Folio to -Folio for some functions and in the realm of -minor-free graphs.
Finally, given positive integers and together with a -colorful graph and a set of size we say that a vertex is irrelevant (or, more precisely, -irrelevant) if the -folio of relative to is equal, up to isomorphism, to the -folio of relative to
7.1 Irrelevant vertices in cliques
In this subsection, we lift the first step from the original algorithm to also work for colorful -Folio.
Let be a -colorful graph and let The -folio of relative to is generic if the -folio of the rooted colorful graph contains every rooted -colorful graph with detail at most where
Let be a rooted -colorful graph and let be disjoint. Let be another rooted -colorful graph. We say that is -restricted if Moreover, we say that is a -extension of if
is -restricted,
and are isomorphic with isomorphism
for all and
Let be a -colorful graph and let Let be a, possibly empty, subset of colors. We say that the -folio of relative to is -generic if
the -folio of the rooted colorful graph contains every -restricted rooted colorful graph with detail at most where and
for every rooted colorful graph that is -restricted and that belongs to the -folio of the rooted colorful graph where any -extension of of detail at most also belongs to the -folio of
The following is an extension of proposition 3.3 with respect to folios.
Proposition 7.1 (Robertson and Seymour [79]).
i)
for every either is connected, or every component of has a vertex in
ii)
iii)
there is no such that there is a separation of of order less than with and
Then the -folio of relative to is generic.
This immediately allows us to deduce a similar statement about colorful rainbow minor models of large cliques.
Corollary 7.2.
Let and be positive integers with Further, let be a -colorful graph and of size
If there exists a colorful minor model of a rainbow in such that there does not exist a separation of order less than such that then the -folio of relative to is generic.
As a final piece, we also need the following extension of Lemma 2.3 from [79]. The proof is straightforward and we omit it here.
Lemma 7.3.
Let and be positive integers. Let be a -colorful graph, and let be a separation in with If a vertex is -irrelevant in then is -irrelevant in
The main theorem of this subsection is now as follows.
Theorem 7.4.
Let and be positive integers with Let be a -colorful graph, with and let be a minor model of in Finally, let be a separation such that
i)
there exists such that
ii)
iii)
there exists a separation of order at most where is maximal and
iv)
has minimum order subject to (i), (ii), and (iii), and
v)
is maximal subject to (i), (ii), (iii), and (iv)
Then for every is -irrelevant.
Proof.
We fix to be a separation meeting the requirements in the assertion. Moreover, let us fix an arbitrary vertex and let and By lemma 7.3 it is sufficient to show that is -irrelevant in To see this it further suffices to show that the -folio of relative to is -generic.
For this, by lemma 3.9, we may assume that there exists a near partition of with and a separation of order at most such that there exists a minor model of in where and Moreover, still by lemma 3.9, we may assume to be connected and, if is the restriction of to then is a colorful minor model of a rainbow in Notice that and thus we may further assume to be maximised by our choice of This implies that
Observe that
Hence, and, moreover, Indeed, if we analyse the situation a bit more carefully we obtain the following inequality:
Since we chose there exists at most one such that If such an exists let us denote it by otherwise let be an arbitrary member of and let Then is a colorful minor model of a rainbow in Moreover, we have that Let be the coloring of obtained from by setting for all By corollary 7.2 we now obtain that the -folio of relative to is generic. Indeed, since for all implies that it follows that the -folio of relative to is -generic as desired. ∎
Robertson and Seymour [79] provide an algorithm that, given a sufficiently large clique-minor model and a set similar to the one in lemma 7.3, finds an irrelevant vertex in time (Algorithm 6.2 in [79]). It follows from the proof of theorem 7.4 that this algorithm is also sufficient to find the irrelevant vertex promised by theorem 7.4 in time
Corollary 7.5.
There exists an algorithm that takes as input a -colorful graph a set of size at most and a minor model of in where and finds a vertex that is -irrelevant in in time where is a computable function.
7.2 Colorful folio checking
In this subsection we now complete our algorithm for Colorful -Folio. To this end, we first present a reduction of the colorful version of folio checking to the non-colorful version in the absence of a fixed minor. This is done by producing an anti-chain of graphs for the minor relation that can then be used to encode colors of vertices.
Let be a function such that, for every even there are pairwise non-isomorphic -regular graphs on vertices.
Lemma 7.6.
For every even integer it holds that
The proof of lemma 7.6 is an interpretation of the asymptotics of some lower bound on given by Liebenau and Wormald [63]. We moved the tedious details of the proof to Appendix B. We also need the following classic result of extremal combinatorics.
Proposition 7.7 ([57, 58, 84]).
There exists a constant such that for every every -minor-free graph is -degenerate.
Lemma 7.8.
There exist a function and a constant such that, for every every -regular graph contains as a minor. Moreover,
Proof.
According to proposition 7.7, every -minor-free graph is -degenerate and therefore it has average degree at most The lemma follows as -regular graphs have average degree ∎
A minor anti-chain is a set of graphs such that no graph in is a minor of another graph in
Lemma 7.9.
There exists a universal constant such that for every there is a minor anti-chain of many connected graphs, each containing as a minor and each of size
Proof.
Let be the set of all pairwise non-isomorphic -regular graphs on vertices. We enhance each graph by adding a new vertex and making it adjacent to all vertices of Let be the collection of all graphs constructed this way. Then every graph in is connected. By a rough estimation, based on the lower bounds in lemma 7.8 and lemma 7.6, there is some universal constant such that The assertion now follows because every graph in has vertices. ∎
Let be a rooted -colorful graph with Given an integer we define the rooted graph where the graph is defined as follows: Let and be a minor anti-chain as guaranteed lemma 7.9. By the choice of we know that contains at least graphs. Let be of the graphs from We start from and, for every vertex of we introduce disjoint and fresh copies of the graphs in and then we introduce edges for each and We refer to the graphs as the -decorations of A decoration of is a -decoration of for some We refer to the vertices of the decorations of as decoration vertices of and we denote the set of decoration vertices by
Considering the construction above, we easily observe the following.
Observation 7.10.
Let be an integer and be a rooted -colorful graph such that does not contain as a minor. Then the graph of excludes as a minor and contains vertices, where
Lemma 7.11.
Let and be rooted -colorful graphs such that both and exclude as a minor, where Then is a colorful minor of if and only if is a minor of
Proof.
For the forward direction let us assume that is a colorful minor model of in For every let be the set of the -decorations of assuming that is isomorphic to By the definition of a colorful minor model, for each color in there is some vertex in where This implies that there is a -decoration of such that and are isomorphic. Now, for every let us define and observe that is a minor model of in Hence, contains as a minor.
For the reverse direction, assume that is a minor model of in where the quantity is minimized.
Again for each vertex and each we denote by the -decoration of that is isomorphic to Also we use for the graph along with the vertex and the edges between and the vertices of As each has at least two vertices, is connected, and is adjacent to every vertex of but no other vertex of has a neighbor in the graph is a block of By lemma 7.9, and therefore as well, contains as a minor. These two facts, the minimization in the choice of and the fact that both and exclude as a minor imply that is a minor of some decoration of for every and Since all possible decorations have the same size and are pairwise non-comparable, this implies that and are isomorphic. This implies that where is the unique vertex of containing the vertices of in its neighborhood. Indeed, this means that for all and This further implies that is a colorful minor model of in as required. ∎
The equivalence of lemma 7.11 along with observation 7.10 imply the following.
Corollary 7.12.
Suppose that there is an algorithm that, given a rooted graph where is -minor-free, computes the -folio of in time for two non-decreasing functions and Then there is an algorithm that, given a rooted -colorful graph where is -minor free, computes the -folio of in time, where
With corollary 7.12 we now have a tool to deduce an algorithm for Colorful -Folio whenever we are in the regime of -minor-free graphs. The remainder of this section is dedicated to piecing everything together to finally deduce Theorems 1.3 and 1.2.
Theorem 7.13.
For all non-negative integers and Colorful -Folio can be solved in time on -colorful graphs on vertices and with edges.
The overall algorithm for theorem 7.13 works as follows. Let be the -colorful graph for which we want to to solve Colorful -Folio.
In a first phase we repeatedly find – until this becomes impossible – a large clique minor in and then use corollary 7.5 find and remove an irrelevant vertex.
Once the first phase is completed, we arrive at a -colorful graph which is an equivalent instance of Colorful -Folio and where excludes a clique-minor of fixed size depending only on and Now we may combine corollary 7.12 with a known algorithm for -Folio to solve the entire problem.
It is important to point out that our estimate of comes from a technical difficulty in the application of corollary 7.12. In order to apply corollary 7.12 we need to be able to assume that or input graph from the -colorful graph already excludes a fixed clique-minor. This introduces the requirement to split the problem into two phases as explained above. In the recent algorithm for the -Folio problem on graphs by Korhonen, Pilipczuk, and Stamoulis [55], both phases happen – roughly – simultaneously, which allows them to achieve a total running time of
Proposition 7.14 (Korhonen, Pilipczuk, Stamoulis [55]).
There exists an algorithm that, given a graph and a set of size at most as input, solves -Folio in time
Indeed, their algorithm implies an almost-linear-time algorithm to check for -minor containment in graphs.
Proposition 7.15 (Korhonen, Pilipczuk, Stamoulis [55]).
There exists an algorithm that, given two graphs and as input, finds a minor-model of in or determines that is -minor-free in time
With this, we have all necessary tools in place to prove theorem 7.13.
Proof of Theorem 7.13.
Let us first fix
We now call proposition 7.15 and ask if contains a -minor. If the answer is yes, then corollary 7.5 will find an irrelevant vertex in time A single application of proposition 7.15 takes time. Moreover, after at most iterations we can be sure that the resulting graph of the newly obtained -colorful graph does not contain a -minor. This completes the first phase in time
With this, we are now in the position for utilizing corollary 7.12. Simply, for every permutation of the vertices in we generate a rooted colorful graph and call the algorithm from proposition 7.14 under the reduction provided by corollary 7.12. In total this takes time and the result is the correct colorful -folio of relative to ∎
To conclude this section observe that both -Colorful Minor Checking and Weak -Colorful -Disjoint Paths are special cases of Colorful -Folio. Hence, Theorems 1.2 and 1.3 are immediate corollaries of theorem 7.13.
For the proof of theorem 1.4, we use the classic approach of by Fellows and Langston [25, 26]. For this, given a colorful minor-monotone -colorful graph parameter for every we define the class and observe that if and only if none of the -colorful graphs in the obstruction set is a colorful minor of Therefore, can be decided by solving instances of the -Colorful Minor Checking problem. Given that, from corollary 2.7, the obstruction set a finite set, we obtain the existence of the algorithm claimed in theorem 1.4.
We emphasize that the arguments underlying the algorithm in theorem 1.4 are non-constructive, as they do not provide a method for explicitly constructing the obstruction set As is already the case for uncolored graphs (i.e., when ), the challenge of developing algorithmic techniques for constructing such sets remains open. Addressing this challenge could lead to constructive – and practically applicable – versions of theorem 1.4 for broad classes of problems on colorful graphs, and thus represents an active direction in current algorithmic research.
@@ 8.1 Star Decompositions of Colorful Graphs
The next two results follow from theorem 5.3 and theorem 3.6, respectively, applied to -colorful graphs. These structural results can be stated as follows:
@@
In the following sections, we present algorithmic applications of corollary 8.2 and corollary 8.3.
@@ 8.2 An Algorithmic Reduction
For every non-negative integer we denote the class of all -colorful graphs by We define a functional problem101010We use the term “functional problem on -colorful graphs” instead of “-colorful graph parameter” to emphasize that we are addressing a computational problem. on -colorful graphs to be a map Our goal is to identify conditions that enable the design of algorithms for solving functional problems on -colorful graphs.
@@
Now consider the star decomposition of where is the vertex set of for each By observation 8.1 we know that Using the algorithm from (i), we compute (which is equal to ) in time, as required. ∎
All subsequent meta-algorithmic results will rely on theorem 8.4.
@@ 8.3 Logics and AMTs on Colorful Graphs
The logic modifies by replacing: with () and with We next define the special quantifier : For a structure and index not in ’s index, if there exists a vertex set such that and, moreover, The quantifier is defined analogously, requiring for the edge set ’s endpoints 111111In the definition of [81] it is demanded that that is the maximum treewidth of an -rooted minor of Clearly, and can be seen as -colorul graph parameters that are equivalent because of the main result in [12]. In this paper we prefer to define using instead of as this is more adequate to the colorful setting of the paper. Also, given a formula we use the notation in order to denote the bidimensionality bound that it imposes on the interpretation of In order to unify notation, in case of a CMSO-formula where there is no restriction on quantification, we assume that
@@
By proposition 8.5 (resp. proposition 8.6), CMSO-definability (resp. -definability) ensures Condition (i) of theorem 8.4 when (resp. ). This naturally raises the question of what assumptions may guarantee Condition (ii). One answer, explored in this section, involves the notion of topological minor folios.
@@ 8.4 Folio Representation
theorem 8.4 has two algorithmic conditions. The first one will be derived from the conditions of propositions 8.5 and 8.6. The second will follow from the ability to replace some part of the input colorful graph with another in which we can route the same set of rooted topological minors (up to a given size, depending on the problem). Provided that this replacement can be performed algorithmically, it will yield an additional combinatorial condition for our algorithmic meta-theorems.
@@
theorem 8.7 associates to every confined boundaried -colorful graph a new one with the same -folio but with a bounded number of vertices. We denote this new confined boundaried -colorful graph by
@@
The algorithm enumerates all confined boundaried -colorful graphs with and returns one with the same -topological-folio as By theorem 8.7, such a graph exists. The computation of can be done in time, according to the main result of [42]. ∎
Folio-Representable Functional Problems.
@@
Given a -colorful graph define as the maximum for which the fusion of contains a -segregated grid as a colorful minor.
8.5 AMTs for colorful graphs
We now have all the necessary ingredients for our AMT’s. We conclude with the following corollary of theorem 8.4 when applied for and taking into account corollary 8.2, proposition 8.5, along with lemma 8.8 and theorem 8.7.
@@ 8.5 AMTs for colorful graphs
Let and be a folio-representable, MSO-definable functional problem (via a formula and a ). There exists an algorithm that, given a -colorful graph computes in time, where
@@
First, apply corollary 8.2 to to obtain a star decomposition of with -width in time. Then apply theorem 8.4: Condition (i) holds by proposition 8.5 where and For Condition ii), we have that is a representation of and the algorithms computing and respectively are given by lemma 8.8 and theorem 8.7 for ∎
If now in the above proof we use corollary 8.3 instead of corollary 8.2, we replace by and we apply proposition 8.6 instead of proposition 8.5 we have the following.
@@
Let and be a folio-representable, -definable functional problem (via formula and a ). There exists an algorithm that, given a -colorful graph computes in time, where
Notice that theorem 8.9 and theorem 8.10 can be viewed as adaptations of proposition 8.5 and proposition 8.6 to optimization problems on colorful graphs, where the combinatorial condition captures how the colored vertices of the input should be distributed in rather than imposing a sparsity condition on the input graph that ignores the colors.
Folio representability for decision problems on colorful graphs
@@
where Recall that in the introduction we presented the decision problem analogues of theorem 8.9 and theorem 8.10, namely theorem 1.9 and theorem 1.10. These can be easily derived as follows: Let and let be a decision problem on -colorful graphs that is folio-representable and CMSO-definable (resp. -definable) by some formula We define a new formula on -colorful graphs so that if and only if the equivalence holds. Then, if and only if We then apply the algorithm of theorem 8.9 (resp. theorem 8.10) to compute the value of If this value is then while if it is then
9 Further research directions
As previously noted, the algorithmic contributions of this paper are limited to showcasing the algorithmic aspects of the colorful minor framework we developed. Below, we outline several promising directions for further algorithmic research on problems involving colorful graphs.
@@ 9.1 Packing and covering colorful minors
As a general research project, one might consider the problem of computing the parameters and for different instantiations of (see section 6.5 for the definitions). Next we give examples of known problems that can be expressed by these parameters, given suitable choices of
@@
First of all, assume that is the -colorful rainbow and observe that by Menger’s theorem, for every -colorful graph it holds that and the two parameters are obviously computable in polynomial time. This is extended by lemma 3.2 implying that, for every positive and are (linearly) equivalent when we consider the -colorful graph defined as follows: is the union of disjoint copies of and, for assigns to the vertex of the colors
Trying to push this further, one may consider a possible duality in the case of four sets and ask whether the following is correct: if there is no --linkage and no --linkage, both of order then there is either a --separator or a --separator of size Such a duality cannot exist, not even approximately, since it corresponds to the -colorful graph where assigns colors to one of the two vertices of and colors to the other one, and this -colorful graph is one of the obstructions for the Erdős Pósa property of -colorful graphs (the counterexample is given by figure 11).
@@
It is easy to observe that both parameters and are colorful minor-monotone, for every instantiation of Therefore, checking whether (resp. ) can be done by an algorithm running in time because of theorem 1.4. However, as already mentioned in the end of section 7.2, this argument is not constructive and only yields the existence of the claimed algorithms.
One may construct an algorithm for checking whether as follows: We first check whether contains a -colorful rainbow using theorem 1.2 in time. If this is the case, then we report a positive answer. If not, then we know that therefore, as is folio representable and -definable, we can apply theorem 1.10 and construct an algorithm running deciding whether in time
On the other hand, a general constructive algorithm for checking whether requires more work. We may certainly begin by checking whether the -colorful rainbow is a colorful minor of and if so, we report negatively. Again, given that we may attempt to apply theorem 1.10 for While is -definable, it is not folio representable, and therefore we cannot apply the replacements used in the proof of theorem 8.4.
However, checking whether is still possible when For this, we may apply the replacement argument from theorem 8.4 by using instead the stronger notion of a -folio of a boundaried graph We denote by - the collection of all -folios of the boundaried graphs derived from after removing vertices. As shown by Fomin, Lokshtanov, Panolan, Saurabh, and Zehavi in [31], one can construct an algorithm that, given outputs, in time, a new boundaried graph such that -- and the graph of excludes a clique as a minor, where the size of the excluded clique depends on a function of and (This follows by repeatedly applying the algorithm from [31, Theorem 6], which identifies irrelevant vertices.)
This enables a replacement argument that produces an equivalent instance whose graph has bounded Hadwiger number. The problem can then be solved in time, again using the decision version of proposition 8.6.
To extend this argument to one would need to prove a “colorful” analogue of [31, Theorem 6], which would extend theorem 7.4 to identify vertices that are irrelevant for every possible hitting set of size While one might reasonably believe such an extension is possible, the details of its construction appear to be quite intricate.

### v2

@@ 7.1 Star Decompositions of Colorful Graphs
The next two results follow from theorem˜5.3 and theorem˜3.6, respectively, applied to -colorful graphs. These structural results can be stated as follows:
@@
In the following sections, we present algorithmic applications of corollary˜7.2 and corollary˜7.3.
@@ 7.2 An algorithmic reduction
For every non-negative integer we denote the class of all -colorful graphs by We define a functional problem111111We use the term “functional problem on -colorful graphs” instead of “-colorful graph parameter” to emphasize that we are addressing a computational problem. on -colorful graphs to be a map Our goal is to identify conditions that enable the design of algorithms for solving functional problems on -colorful graphs.
@@
Now consider the star decomposition of where is the vertex set of for each By observation˜7.1 we know that Using the algorithm from (i), we compute (which is equal to ) in time, as required. ∎
All subsequent meta-algorithmic results will rely on theorem˜7.4.
@@ 7.3 Logics and AMTs on colorful graphs
The logic modifies by replacing: with () and with We next define the special quantifier : For a structure and index not in ’s index, if there exists a vertex set such that and, moreover, The quantifier is defined analogously, requiring for the edge set ’s endpoints 121212In the definition of [81] it is demanded that that is the maximum treewidth of an -rooted minor of Clearly, and can be seen as -colorful graph parameters that are equivalent because of the main result in [12]. In this paper we prefer to define using instead of as this is more adequate to the colorful setting of the paper. Also, given a formula we use the notation in order to denote the bidimensionality bound that it imposes on the interpretation of In order to unify notation, in case of a CMSO-formula where there is no restriction on quantification, we assume that
@@
By proposition˜7.5 (resp. proposition˜7.6), CMSO-definability (resp. -definability) ensures Condition (i) of theorem˜7.4 when (resp. ). This naturally raises the question of what assumptions may guarantee Condition (ii). One answer, explored in this section, involves the notion of topological minor folios.
@@ 7.4 Folio representation
theorem˜7.4 has two algorithmic conditions. The first one will be derived from the conditions of propositions˜7.5 and 7.6. The second will follow from the ability to replace some part of the input colorful graph with another in which we can route the same set of rooted topological minors (up to a given size, depending on the problem). Provided that this replacement can be performed algorithmically, it will yield an additional combinatorial condition for our algorithmic meta-theorems.
@@
theorem˜7.7 associates to every confined boundaried -colorful graph a new one with the same -folio but with a bounded number of vertices. We denote this new confined boundaried -colorful graph by
@@
The algorithm enumerates all confined boundaried -colorful graphs with and returns one with the same -topological-folio as By theorem˜7.7, such a graph exists. The computation of can be done in time, according to the main result of [42]. ∎
7.5 AMTs for colorful graphs
We now have all the necessary ingredients for our AMT’s. We conclude with the following corollary of theorem˜7.4 when applied for and taking into account corollary˜7.2, proposition˜7.5, along with lemma˜7.8 and theorem˜7.7.
@@ 7.5 AMTs for colorful graphs
Let and be a folio-representable, MSO-definable functional problem (via a formula and a ). There exists an algorithm that, given a -colorful graph computes in time, where is the maximum monodimensionality of a color in .
@@
Let be the -colorful graph such that for each vertex if and only if and is the color of maximum monodimensionality in Note that the monodimensionality of in is of order (provided that is a fixed number).
First, apply corollary˜7.2 to to obtain a star decomposition of with -width in time. Then apply theorem˜7.4: Condition (i) holds by proposition˜7.5 where and For Condition ii), we have that is a representation of and the algorithms computing and respectively are given by lemma˜7.8 and theorem˜7.7 for ∎
If now in the above proof we use corollary˜7.3 instead of corollary˜7.2, we replace by and we apply proposition˜7.6 instead of proposition˜7.5 we have the following.
@@
Let and be a folio-representable, -definable functional problem (via formula and a ). There exists an algorithm that, given a -colorful graph computes in time, where is the maximum Hadwiger number of a color in .
Notice that theorem˜7.9 and theorem˜7.10 can be viewed as adaptations of proposition˜7.5 and proposition˜7.6 to optimization problems on colorful graphs, where the combinatorial condition captures how the colored vertices of the input should be distributed in rather than imposing a sparsity condition on the input graph that ignores the colors.
Folio representability for decision problems on colorful graphs.
@@
where Recall that in the introduction we presented the decision problem analogues of theorem˜7.9 and theorem˜7.10, namely theorem˜1.8 and theorem˜1.9. These can be easily derived as follows: Let and let be a decision problem on -colorful graphs that is folio-representable and CMSO-definable (resp. -definable) by some formula We define a new formula on -colorful graphs so that if and only if the equivalence holds. Then, if and only if We then apply the algorithm of theorem˜7.9 (resp. theorem˜7.10) to compute the value of If this value is then while if it is then
8 Conclusion
As previously noted, the algorithmic contributions of this paper are limited to showcasing the algorithmic aspects of the colorful minor framework we developed. Below, we outline several promising directions for further research on problems involving colorful graphs.
@@ 8.1 Packing and covering colorful minors
As a general research project, one might consider the problem of computing the parameters and for different instantiations of (see section˜6.5 for the definitions). Next, we give examples of known problems that can be expressed by these parameters, given suitable choices of
@@
First of all, assume that is the -colorful rainbow and observe that by Menger’s theorem, for every -colorful graph it holds that and the two parameters are obviously computable in polynomial time. This is extended by lemma˜3.2 implying that, for every positive and are (linearly) equivalent when we consider the -colorful graph defined as follows: is the union of disjoint copies of and, for assigns to the vertex of the colors
Trying to push this further, one may consider a possible duality in the case of four sets and ask whether the following is correct: if there is no --linkage and no --linkage, both of order then there is either a --separator or a --separator of size Such a duality cannot exist, not even approximately, since it corresponds to the -colorful graph where assigns colors to one of the two vertices of and colors to the other one, and this -colorful graph is one of the obstructions for the Erdős Pósa property of -colorful graphs (the counterexample is given by figure˜11).
@@
It is easy to observe that both parameters and are colorful minor-monotone, for every instantiation of Therefore, checking whether (resp. ) can be done by an algorithm running in time because of theorem˜1.7. However, as already commented after theorem˜1.7, this argument is not constructive and only yields the existence of the claimed algorithms.
One may construct an algorithm for checking whether as follows: We first check whether contains a -colorful rainbow using theorem˜1.6 in time. If this is the case, then we report a positive answer. If not, then we know that therefore, as is folio representable and -definable, we can apply theorem˜1.9 and construct an algorithm running deciding whether in time
On the other hand, a general constructive algorithm for checking whether requires more work. We may certainly begin by checking whether the -colorful rainbow is a colorful minor of and if so, we report negatively. Again, given that we may attempt to apply theorem˜1.9 for While is -definable, it is not folio representable, and therefore we cannot apply the replacements used in the proof of theorem˜7.4.
However, checking whether is still possible when For this, we may apply the replacement argument from theorem˜7.4 by using instead the stronger notion of a -folio of a boundaried graph We denote by - the collection of all -folios of the boundaried graphs derived from after removing vertices. As shown by Fomin, Lokshtanov, Panolan, Saurabh, and Zehavi in [32], one can construct an algorithm that, given outputs, in time, a new boundaried graph such that -- and the graph of excludes a clique as a minor, where the size of the excluded clique depends on a function of and (This follows by repeatedly applying the algorithm from [32, Theorem 6], which identifies irrelevant vertices.)
This enables a replacement argument that produces an equivalent instance whose graph has bounded Hadwiger number. The problem can then be solved in time, again using the decision version of proposition˜7.6.
To extend this argument to one would need to prove a “colorful” analogue of [32, Theorem 6], which would be able to identify a vertex that is irrelevant for every possible hitting set of size While one might reasonably believe such an extension is possible, the details of its construction appear to be quite intricate.
@@
8.2 Extending graph parameters to colorful ones
Recently, two main mechanisms have been proposed to extend treewidth to annotated graphs (i.e., -colorful graphs). These are the rainbow extension [84] and the torso extension [46, 45]. We propose here an extension of both mechanisms to -colorful graphs that applies to every minor-monotone parameter.
Let and let be a minor-monotone graph parameter. We define the torso -extension of and the rainbow -extension of as the -colorful graph parameters and , where
As mentioned in the introduction, , the torso treewidth of an annotated graph, has already been proposed—using different terminology—by Jansen and Swennenhuis in [46] and by Hodor, La, Micek, and Rambaud in [45]. Also, it is the the main combinatorial engine in the proof of theorem˜7.9.
On the other hand, , the rainbow treewidth of an annotated graph, is equivalent to the bidimensionality , which is , where denotes the maximum for which contains the -grid as a minor. This leads us to view (which, again, is equivalent to ) as the correct notion of bidimensionality for colorful graphs. In this sense, measures the “simultaneous presence” of all colors across the bidimensional surface of a grid.
Recall that, as already proved in [45], is equivalent to the maximum for which contains the -segregated grid as a colorful minor. Moreover, by theorem˜1.3, this equivalence extends to colorful graphs with any number of colors. That is, is equivalent to the maximum for which contains one of the parametrically-distinct -segregated grids as a colorful minor (see section˜6.3). This motivates us to interpret this quantity as the monodimensionality of a colorful graph, as it measures the “segregated presence” of all colors along the monodimensional boundary of a grid.
A natural question is when the two aforementioned mechanisms for generating colorful extensions yield equivalent parameters. Let be a minor-monotone graph parameter. It is easy to see that for every -colorful graph , we have , regardless of the choice of or . However, it is not generally the case that there exists a function such that for all . Thus, the question becomes:
For which minor-monotone parameters does it hold that and are equivalent?
We say that a graph parameter is unbounded in a class if the set is infinite. For two graph parameters and , we write if there is a function such that, for every graph , . We conjecture the following.
Conjecture 8.1.
Let and let be a minor-monotone parameter, where . Then
if , then and are equivalent if and only if is unbounded in outerplanar graphs;
if , then and are equivalent if and only if is unbounded in paths;
if , then and are not equivalent.
Proof of the “only if” directions of conjecture˜8.1.
We start with the case where . We use for the -segregated grid and we assume that is bounded on outerplanar graphs. Observe first that, for every , all rainbow minors of are outerplanar, therefore the value of on a is upper bounded by some constant that is independent of the value of . On the other side, we claim that the value of on is lower bounded by some (increasing) function of . To prove this claim, assume to the contrary, that there is some constant , such that for every , . This means that there is a where . But as , this means that is also always upper bounded by some constant . Notice now that is a -connected set of , i.e., and for all subsets with there are disjoint - paths in . Because , is a -connected set also in , which implies that is lower bounded by for some constant (see [20, Proposition 3]). This means that . We now have a contradiction by picking to be big enough so that and the claim holds. We just proved that the values of and may differ arbitrarily on . Therefore and we may conclude that and are not equivalent colorful graph parameters.
The proof for goes in a similar way. We use for the -segregated grid and we assume that . The fact that is bounded by a constant, no matter the choice of , follows by the fact that every rainbow minor of is the disjoint union of paths. Assume now that there is some , that does not depend on , and a set where (and therefore also ) and, for every component of , it holds that either or . Using this and the fact that there are vertex disjoint paths in between and , one can easily conclude that, provided that , for every component of , contains at most vertices from . This in turn implies that , where . The proof now goes as in the case where because of the fact that is a -connected set of which, for hight enough values of , contradicts the fact that .
For the case , notice that every rainbow parameter is bounded in any -segregated grid as these graphs do not contain rainbow minors. The fact that the torso extension of is unbounded on such graphs is a straightforward extension of the proof when , again using the fact that for every two distinct colors , there are vertex disjoint - paths. ∎
A proof of the “if” direction of conjecture˜8.1 (if exists) seems to be more involved and we leave it as an open problem.
A typical example where conjecture˜8.1 holds is pathwidth,131313We do not give the definitions of the parameters pathwidth and treedepth here. For our purposes, it is enough to mention that a graph class has bounded pathwidth/treedepth if and only if it does not contain all trees/paths. denoted by . While and are equivalent when (because of the results of [45]), this equivalence fails when because pathwidth is bounded in paths. For the case , notice that outperplanar graphs have unbounded pathwidth.
Another example is treedepth, denoted by . In this case, it is possible to verify that and are equivalent for both (for this one may apply theorem˜1.3 in order to bound torso treewidth and then recycle the proof of [67, Theorem 12.12]). For the cases , , notice that paths, as well as outperplanar graphs, have unbounded treedepth.